# FusionDTI: Fine-grained Binding Discovery with Token-level Fusion for Drug-Target Interaction

### v1

@@ 4.8 Case Study
A further strength of FusionDTI to enable explainability, which is critical for drug design efforts, is the visualisation of each token’s contribution to the final prediction through cross-attention maps. To compare with the DrugBAN model, we examine three identical pairs of DTI from the Protein Data Bank (PDB) [4]: (EZL - 6QL2 [17], 9YA - 5W8L [28] and EJ4 - 4N6H [12]). As shown in Table 4, our proposed model predicts additional binding sites (in bold) evidenced by PDB [4] in comparison to the DrugBAN model. For instance, to predict the interaction of the drug EZL with the target 6QL2, our proposed model using BertViz [38] highlights potential binding sites as illustrated in Figure 8. Our CAN module is effective in capturing fine-grained binding information at the token level. In particular, we address the lack of structural information on protein sequences by employing the SA vocabulary, which matches each residue to a corresponding 3D feature via Foldseek [35]. This study highlights the effectiveness of FusionDTI in enhancing performance on the DTI task, thereby supporting more targeted and efficient drug development efforts. In Section 6.5 of the Appendix, we present three pairs of prediction visualisations.
@@ 5 Conclusions
With the rapid increase of new diseases and the urgent need for innovative drugs, it is critical to capture and gauge fine-grained interactions, since the binding of specific drug atoms to the main amino acids is key to the DTI task. Despite some achievements, fine-grained interaction information is not effectively captured. To address this challenge, we introduce FusionDTI uses token-level fusion to effectively obtain fine-grained interaction information between drugs and targets.

### v2

@@ 4.8 Case Study
A further strength of FusionDTI to enable explainability, which is critical for drug design efforts, is the visualisation of each token’s contribution to the final prediction through cross-attention maps. To compare with the DrugBAN model, we examine three identical pairs of DTI from the Protein Data Bank (PDB) [4]: (EZL - 6QL2 [18], 9YA - 5W8L [29] and EJ4 - 4N6H [13]). As shown in Table 5, our proposed model predicts additional binding sites (in bold) evidenced by PDB [4] in comparison to the DrugBAN model. For instance, to predict the interaction of the drug EZL with the target 6QL2, our proposed model using BertViz [39] highlights potential binding sites as illustrated in Figure 8. Specifically, our CAN module is effective in capturing fine-grained binding information at the token level, as we have successfully predicted the novel binding between Gln92 and the benzothiazole ring [11]. In particular, we address the lack of structural information on protein sequences by employing the SA vocabulary, which matches each residue to a corresponding 3D feature via Foldseek [36]. This study highlights the effectiveness of FusionDTI in enhancing performance on the DTI task, thereby supporting more targeted and efficient drug development efforts. In Appendix 6.5, we present three pairs of prediction visualisations.
@@ 5 Conclusions
With the rapid increase of new diseases and the urgent need for innovative drugs, it is critical to capture and gauge fine-grained interactions, since the binding of specific drug atoms to the main amino acids is key to the DTI task. Despite some achievements, fine-grained interaction information is not effectively captured. To address this challenge, we introduce FusionDTI uses token-level fusion to effectively obtain fine-grained interaction information between drugs and targets. Through experiments on three well-known benchmark datasets, we demonstrate that our proposed FusionDTI model outperforms eight state-of-the-art baselines in DTI prediction, particularly in the more realistic cross-domain scenario. Additionally, we show that the attention weights of the token-level fusion module in our model can highlight potential binding sites, providing a certain level of explainability.